OpenStudy (anonymous):

(3/7)2 dont know how to draw this

OpenStudy (alexandervonhumboldt2):

3/7*2=(3*2)/7

OpenStudy (alexandervonhumboldt2):

6/7

OpenStudy (alexandervonhumboldt2):

or if 3/7^2=9/49

OpenStudy (anonymous):

thanks i need all the help i can get